# Text-centric Alignment for Multi-Modality Learning

### v1

@@ Abstract.
This research paper addresses the challenge of modality mismatch in multimodal learning, where the modalities available during inference differ from those available at training. We propose the Text-centric Alignment for Multi-Modality Learning (TAMML) approach, an innovative method that utilizes Large Language Models (LLMs) with in-context learning and foundation models to enhance the generalizability of multimodal systems under these conditions. By leveraging the unique properties of text as a unified semantic space, TAMML demonstrates significant improvements in handling unseen, diverse, and unpredictable modality combinations. TAMML not only adapts to varying modalities but also maintains robust performance, showcasing the potential of foundation models in overcoming the limitations of traditional fixed-modality frameworks in embedding representations. This study contributes to the field by offering a flexible, effective solution for real-world applications where modality availability is dynamic and uncertain.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06copyright: none
@@ 1. Introduction
This work considers a foundation model for multimodal learning. Multimodal learning focuses on integrating and interpreting information across various data modalities during a model’s training and inference stages. This interdisciplinary field intersects with domains such as computer vision, natural language processing, audio signal processing and programming (Xu et al., 2024; Zhang et al., 2024; Tsai et al., 2023; Song et al., 2023). The primary goal is to enable machine learning models to comprehend and integrate different forms of data, such as combining text, images, and audio into a cohesive understanding. The challenge in multimodal learning lies in effectively capturing and integrating the nuanced and often complex relationships between different data types. This requires recognizing patterns within each modality and understanding how these patterns interact and complement each other across modalities. The applications of multimodal learning are diverse and far-reaching, enhancing accessibility through integrated speech, text, and visual information services and contributing to more immersive experiences in areas like virtual reality and interactive systems. As the field evolves, it holds the potential to revolutionize human-computer interaction, content creation, and a multitude of other sectors by enabling a more sophisticated and intuitive understanding of complex, multimodal data.
Figure 1. Recent works for cross-modality learning focus on different modality mismatch scenarios. This paper aims to establish a general method for all mismatch types and combinations. Specifically, our experiments focus on generalizability that could be transferred to an unseen modality at test time.
The motivation for this research arises from the dynamic nature of real-world data, where modalities can unpredictably vary or be absent at inference time. Traditional multimodal learning methods, typically fixated on static modality combinations during both training and inference, fall short in such fluid environments. Although foundation models pre-trained on vast datasets across numerous modalities are frequently employed to encode these modalities into embeddings, changes in the input modality necessitate retraining the discriminative downstream model for accurate predictions. For instance, a classifier trained on embeddings from images and tabular data might not be suitable for inferring audio inputs due to disparities in embedding dimensions and semantic content. To overcome these challenges, our research leverages text’s capacity to serve as a unified semantic space alongside the extensive zero-shot prediction capabilities of Large Language Models (LLMs). Text’s modality-invariant nature provides a versatile bridge across different data types, potentially circumventing issues like modality collapse and extending generalizability to unseen modalities. Moreover, advanced text analysis tasks such as translation, summarization, and explanation have been extensively researched and integrated into Large Language Models (LLMs). This integration offers powerful capabilities to align various modalities effectively. By employing foundation models to convert diverse modalities into text, we propose a novel and flexible multimodal learning framework, offering a promising solution to the limitations of current multimodal learning methodologies for unseen modalities.
@@
Our research objective is to develop a downstream model that is invariant across modalities, capable of being trained on data of certain modalities and performing zero-shot predictions on unseen modalities at test time. For instance, assuming only image, tabular, and text data are available during training, we want to train a model that can take video inputs during inference time, as shown in Figure 3. Note that, as we will demonstrate through experiments, simply converting all modalities into text for training and inference is insufficient, as the mismatch between different modalities does not translate and align as seamlessly in text, especially in complex scenarios involving multiple modality combinations. To address these challenges, our work has extensively explored challenges such as modality alignment, translation, missing modality imputation, and generation. Our goal is to pioneer methods in multimodal learning that leverage text representations and in-context learning, thereby establishing a seamless integration across various modalities. Furthermore, we aim to explore the balance between flexibility and performance when employing text representations for multimodal learning, in contrast to traditional embedding approaches.
@@
TAMML employs Large Language Models (LLMs) for data transformation across various modalities, with the aim of creating a unified semantic space. This process is conducted exclusively through in-context learning. Initially, we transform different modalities into text. Recently, various solutions have been developed, such as GPT-4 (OpenAI, 2023), Blip2 (Li et al., 2023) for vision, and TabLLM (Hegselmann et al., 2023) for tabular data. Following this, we engage LLMs in text-style translation across modalities, ensuring that all modalities in their textual representation adopt a consistent linguistic structure, thereby reducing the gap between different modalities. To further align these modalities within a closer semantic space, remove redundant information, and mitigate the heterogeneity inherent in text data from diverse sources, we further conduct modality summarization. This step involves a concise summarization of the translated data. Additionally, TAMML includes a reasoning augmentation step akin to the Chain-of-Thought (Wei et al., 2022) method, where we enhance the data with LLMs to boost prediction and judgment capabilities. Moreover, we leverage LLMs as a source of large-scale external knowledge, enriching the data understanding and interpretative depth (Chen et al., 2023). We would like to answer several hypotheses through extensive experiments. First, whether TAMML is more effective compared with the existing solutions in predicting data of unseen modality. Second, although this work mainly focuses on the prediction given unseen modality, we would like to understand whether the proposed solution is still effective when the modality in testing is already seen during training. Third, whether the text-as-the-medium strategy is generally more robust compared to the solutions focusing on embedding-based cross-modality transfer. We benchmarked TAMML against existing methodologies in closely related tasks, particularly focusing on zero-shot learning cross-modality translation, which involves translating unseen source data to a different target domain. Techniques like MIDiffusion (Wang et al., 2023) and SDEdit (Meng et al., 2022) demonstrate commendable performance in tasks such as domain translation within images. However, these methods encounter challenges when the source and target domains represent completely different modalities.
@@
  • We demonstrate TAMML can significantly outperform SOTA approaches by conducting multiple experiments on real-world datasets. We also have an ablation study to analyze the effectiveness of each component in TAMML .
@@ 2. Related Works
In this section, we begin by briefly exploring the advancements of multimodal foundation models in Section 2.1. Next, we discuss zero-shot learning cross-modality translation in Section 2.2.
@@ 2.1. Multimodal Foundation Models
Extending beyond single-modality applications, large multimodal language models (MLLMs) have shown remarkable proficiency in both reasoning and generation tasks (Yin et al., 2023). For instance, Flamingo (Alayrac et al., 2022) employs a vision encoder for feature extraction from images and utilizes a resampler to reduce computational complexity. This is achieved by learning a predefined number of latent input queries, which are then processed by a Transformer that cross-attends to the visual features. On the other hand, Kosmos-2 (Peng et al., 2023) adopts a more streamlined architecture, comprising only a vision encoder, a linear projector, and a base transformer. This model directly concatenates image and text features, using unique tokens for differentiation. LLaVA (Liu et al., 2023) represents a novel end-to-end trained large multimodal model that combines a vision encoder and Vicuna (Chiang et al., 2023) for general-purpose visual and language understanding, achieving impressive chat capabilities mimicking spirits of the multimodal GPT-4. However, despite the impressive capabilities of MLLMs in handling multimodal data, acquiring large-scale datasets encompassing multiple modalities remains a significant challenge.
Figure 3. In the training phase, each modality in raw input is first transformed into text representations using a corresponding foundation model. For example, we employ an image caption model for images and TabLLM (Hegselmann et al., 2023) to transform tabular data. Following the modality transformation, we utilize LLM for text-style translation across modalities, applying a one-shot in-context learning approach to adapt the linguistic style as anticipated during inference. After text-style translation, modality summarization and reasoning augmentation are applied in parallel. Finally, the output texts are concatenated as the training inputs to a transformer model for downstream prediction. The inference phase follows a similar pattern, with the notable exception of bypassing the translation module to preserve the original linguistic style.
@@ 3. Methodologies
This section focuses on how TAMML makes training modality adapt to unseen testing modality and unseen modality combinations. In Section 3.1, we define the problem scenario and its notation of our research, and Section 3.2 describes the individual text transformation method to different modalities. Next, we introduce our main algorithm to connect information from multiple modalities. Section 3.3 translating training modalities to inference modalities. Section 3.4 summarizes different modalities into human-written descriptions, and Section 3.5 leverages external knowledge as data augmentation. At last, Section 3.6 elucidates the flow of training and inference within the downstream task.
@@ 3.1. Problem Formalization
Suppose we have a set $M$ of $p$ modalities $M=\{m_{1},m_{2},\ldots,m_{p}\}$. In the training phase, a subset of modalities $M_{T}\subseteq M$ is used. In the inference phase, a different subset $M_{I}\subseteq M$ is utilized, with the critical condition that $M_{T}\cap M_{I}=\emptyset$ to ensure no overlap in modalities between training and inference.
Within this framework, we define two distinct datasets for the training and inference phase. The training dataset $D_{T}$ is composed of a set of $n_{T}$ samples, and each sample $x$ is restricted to $M_{T}$, denoted as $D_{T}=\{(x_{M_{T}}^{i},y^{i})\}_{i=1}^{n_{T}}$. Similarly, the inference dataset consists of a set of $n_{I}$ samples restricted to $M_{I}$, formalized as $D_{I}=\{(x_{M_{I}}^{i},y^{i})\}_{i=1}^{n_{I}}$. Our algorithms are designed to build the model $F$ on $D_{T}$ and evaluate unseen data with unseen modality combinations in $D_{I}$ to measure the model’s ability to generalize knowledge in zero-shot multimodal learning. Besides, we design a dataset $D_{E}$ containing $k$ samples, each of them is restricted to $M$. This dataset is used when we prompt.
@@ 3.2. Text Transformation
For the challenge of the heterogeneity between different modalities, we attempt to map the data of different modalities into a similar semantic space at the raw input stage, integrating modalities with minimal processing.
For image modality, we utilize the state-of-the-art image caption model, which allows us to take images and answer questions about them. In this stage, the image caption model is prompted to describe images and generate detailed text output, effectively translating visual content into descriptive text. For table modality, we implement the simple text serialization template, which is the textual enumeration of all features as ”The column name values,” proposed by TabLLM(Hegselmann et al., 2023). According to TabLLM’s study, this template is proven to outperform zero-shot tabular learning in the LLM. For text modality, we maintain human-written sentence forms without conversion to maintain their original natural linguistic structure. After transformation, the sentence order of each sample $x$ is determined by the order of modalities in $M$ to concatenate the text sentences and is used as the input for the next step. We have included and compared various SOTA image caption models and LLMs in Appendix C.2.
@@ 3.3. Text-style Translation across Modality
After text transformation, we obtain textual representation for each modality. Next, we employ Large Language Models (LLMs) to perform text-style translation across different modalities. This is because, although all types of data are converted into textual representation, there are still syntactic and semantic gaps between the transformed text across different modalities, which are highly dependent on the foundation models used for transformation. For instance, the textual representation of tabular data can follow a static pattern (e.g., ”the type of the pet is Cat; the name of the pet is Kali,” etc.), while the textual representation of the image is affected by the language style of the image caption model (e.g., ”the image shows that a young kitty with a coat…”). This divergence in textual representation leads to gaps in semantic space distribution. Our solution is to adopt in-context learning in LLM to translate the text style to the one in the target space, reducing the mismatch between training and inference data. Text-style Translation function $T$ can be formulated as follows:
$T:M_{T}\rightarrow M_{I}$
Specifically, we enhance the capability of the LLM to effectively interpret the semantic content and convert the modality text from $M_{T}$ format into $M_{I}$ format. This transformation is achieved through few-shot in-context learning, where LLM is exposed to three examples, which are samples from $D_{E}$, illustrating the desired translation within the context of our modality combination for training and inference purposes. Moreover, to mitigate information loss during conversion, we specifically emphasize that the integrity of the original information should be maintained in our prompt. This emphasis ensures that the essential information content and cross-modal connections are preserved during the cross-modal translation process. Appendix A Fig. 5 demonstrates the details of the Text-style Translation prompt.
To achieve the two effects of reduction of the distance gap between modalities and information preservation, the LLM needs to achieve $T$ is optimized to minimize the semantic distance $d$ between the $D_{T}$ and $D_{I}$, while preserving the core information content $I_{c}$. Thus, the LLM optimization objective can be formulated as:
$\min_{T}d(T(D_{T}),D_{I})\quad\text{subject to}\quad I_{c}(T(D_{T}))=I_{c}(D_{T})$

### v2

# Text-centric Alignment for Multi-Modality Supervised Learning

@@ Abstract.
This paper addresses the challenge of modality mismatch in supervised learning, where the modalities available during inference differ from those available during training. We propose an innovative method that utilizes Large Language Models with in-context learning and foundation models to enhance the generalizability of multimodal systems under these conditions. By leveraging the unique properties of text as a unified semantic space, this paper demonstrates significant improvements in handling unseen, diverse, and unpredictable modality combinations. The proposed solution not only adapts to varying modalities but also maintains robust performance, showcasing the potential of foundation models in overcoming the limitations of traditional fixed-modality frameworks in embedding representations. This study contributes to the field by offering a flexible, effective solution for real-world applications where modality availability is dynamic and uncertain.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; October 21–25, 2024; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06copyright: none
@@ 1. Introduction
Figure 1. This paper establishes a general method for all mismatch types and combinations. As the figure shows, the model utilizes three modalities during training. Our unified model handles inference for any combination of modalities, such as User1’s unseen audio-video combination or the diverse combinations presented by User2 and User3.
This work targets the challenge of developing a supervised learning model where the input modality in the testing (or prediction) phase differs from that in training. The motivation for this research arises from the dynamic nature of real-world data, where modalities can unpredictably vary or even be absent at inference time. Consider the following scenarios: 1. A hospital has extensive image and text data about its patients, such as X-ray images and doctors’ written diagnoses. This data can be used to train an AI model that diagnoses patients based on both image and text inputs. However, to enhance patient satisfaction, the hospital wants to develop a dialog system that can diagnose patients based on their audio descriptions of their symptoms. Typically, achieving this would require collecting audio data and employing transfer learning or domain adaptation techniques to align the information across modalities, a process that demands significant additional effort, cost, and time. 2. In the financial industry, a bank aims to train an AI assistant to recommend suitable financial products to customers. The existing data comprises millions of customers’ past credit card transactions and purchase histories in the form of time series and tables. However, during the inference phase, the AI assistant must interact with customers using only text messages. Collecting sufficient text data to build a model to align text with time-series and transaction records can be expensive and resource-intensive.
Traditional multimodal learning methods, typically fixated on static modality combinations during both training and inference, fall short in such fluid environments. Therefore, this paper explores the possibility of creating a supervised model that utilizes only existing modalities (e.g., images and text in the hospital scenario, and time series and tables in the bank scenario) during training, yet allows for the incorporation of an unseen modality (e.g., audio signals or text inputs) during inference. If successful, this approach would eliminate the need for additional data collection and modality alignment, thereby reducing the associated costs and efforts.
There are two plausible directions to tackle this challenge.
First, we can rely on universal models pre-trained on vast datasets across numerous modalities to encode these modalities into embeddings. However, changes in the input modality necessitate retraining the discriminative downstream model for accurate predictions. From our experiments, we infer a classifier trained on embeddings from images and tabular data is not ideal for making inferences about audio inputs due to disparities in embedding dimensions and semantic content.
The second direction, which this paper adopts, is to convert every modality into a single modality and build the model based on that unified modality. We argue that converting all modalities into text could be a favorable choice. Text can serve as a unified semantic space, leveraging the extensive zero-shot prediction capabilities of Large Language Models (LLMs). The modality-invariant nature of text provides a versatile bridge across different data types, potentially circumventing issues like modality collapse and extending generalizability to unseen modalities. Furthermore, advanced text analysis tasks such as translation, summarization, and explanation have been extensively researched and integrated into LLMs, offering powerful capabilities to align various modalities effectively.
@@
Our objective is to develop a downstream model that is invariant across modalities. This model should be capable of being trained on data of certain modalities and performing zero-shot predictions on various modality combinations during testing, regardless of whether they were seen during training, as shown in Figure 1 and Figure 3. Note that, as we will demonstrate through experiments, simply converting all modalities into text for training and inference is insufficient, as the mismatch between different modalities does not translate and align as seamlessly in text, especially in complex scenarios involving multiple modality combinations. To address these issues, our work has explored challenges such as modality alignment, translation, and augmentation. Our goal is to pioneer methods in multimodal learning that leverage text representations and in-context learning, establishing seamless integration across various modalities. Furthermore, we aim to explore the balance between flexibility and performance when employing text representations for multimodal learning, in contrast to traditional embedding approaches.
@@
TAMML employs LLMs for data transformation across various modalities, with the aim of creating a unified semantic space. This process is conducted exclusively through in-context learning. Initially, we transform different modalities into text. Recently, various solutions have been developed, such as GPT-4 (OpenAI, 2023), Blip2 (Li et al., 2023b) for vision, and TabLLM (Hegselmann et al., 2023) for tabular data. Following this, we engage LLMs in text-style translation across modalities, ensuring that all modalities in their textual representation adopt a consistent linguistic structure, thereby reducing the gap between different modalities. To further align these modalities within a closer semantic space, remove redundant information, and mitigate the heterogeneity inherent in text data from diverse sources, we further conduct modality summarization. This step involves a concise summarization of the translated data. Additionally, TAMML includes a reasoning augmentation step akin to the Chain-of-Thought (Wei et al., 2022) method, where we enhance the data with LLMs to boost prediction and judgment capabilities. Moreover, we leverage LLMs as a source of large-scale external knowledge, enriching the data understanding and interpretative depth (Chen et al., 2023). We aim to answer several hypotheses through extensive experiments. First, whether TAMML is more effective compared to existing solutions in predicting data of unseen modalities. Second, although this work focuses on predicting unseen modalities, we want to understand whether the proposed solution is effective when the modality in testing is already seen during training. Third, whether the text-as-the-medium strategy is more robust compared to embedding-based cross-modality transfer solutions. We benchmarked TAMML against existing methodologies in closely related tasks, particularly focusing on zero-shot learning cross-modality translation, which involves translating unseen source data to a different target domain. Techniques like MIDiffusion (Wang et al., 2023a) and SDEdit (Meng et al., 2022) demonstrate commendable performance in tasks such as domain translation within images. However, these methods encounter challenges when the source and target domains represent completely different modalities.
@@
  • We demonstrate that TAMML can significantly outperform SOTA approaches by conducting multiple experiments on real-world datasets. We also have an ablation study to analyze the effectiveness of each component in TAMML .
@@ 2.1. Multimodal Foundation Models
Extending beyond single-modality applications, large multimodal language models (MLLMs) have shown remarkable proficiency in both reasoning and generation tasks (Yin et al., 2023). For instance, Flamingo (Alayrac et al., 2022) employs a vision encoder for feature extraction from images on top of the transformer backbone. On the other hand, Kosmos-2 (Peng et al., 2023) is comprised of a vision encoder, a linear projector, and a base transformer. These models directly concatenate image and text features, using unique tokens for differentiation. However, despite the impressive capabilities of MLLMs in handling multimodal data, acquiring large-scale datasets encompassing multiple modalities remains a significant challenge.
Several works involve processing images with foundation models and combining the results into text for LLM training, linking visual information with text. For example, LLaVA (Liu et al., 2023) utilizes GPT4 to transfer images into captions and object detection results into text descriptions as their multimodal training data. VideoChat-Text (Li et al., 2023a) encodes video into textual descriptions. In medicine, OphGLM (Gao et al., 2023) uses classification and segmentation models to extract information from fundus images and create diagnostic reports for LLMs. Similarly, ChatCAD (Wang et al., 2023b) transforms X-ray outputs from CAD models into natural language for LLM input.
Figure 3. In the training phase, each raw input modality is transformed into text representations using a corresponding foundation model. Following the modality transformation, summarization, and augmentation are applied in parallel. Finally, the output texts are concatenated as the training inputs to a transformer model for downstream prediction. The inference phase follows a similar pattern, with the exception of utilizing an LLM for the text-style translation after the text transformation module. We apply a one-shot in-context learning approach to adapt the linguistic style as anticipated during training.
@@ 3. Methodologies
This section focuses on how TAMML enables training modalities to adapt to unseen testing modalities and unseen modality combinations. In Section 3.1, we define the problem scenario and its notations, and Section 3.2 describes the individual text transformation methods for different modalities. Next, we introduce our main algorithm to connect information from multiple modalities. Section 3.3 discusses translating inference modalities to training modalities. Section 3.4 summarizes different modalities into human-written descriptions, and Section 3.5 leverages external knowledge as data augmentation. Lastly, Section 3.6 elucidates the flow of training and inference within the downstream task. The whole process is shown in Figure. 3.
@@ 3.1. Problem Formalization
Suppose we have a set $M$ of $p$ modalities, $M=\{m_{1},m_{2},\ldots,m_{p}\}$. In the training phase, a subset of modalities $M_{T}\subseteq M$ is used. In the inference phase, a different subset $M_{I}\subseteq M$ is utilized. This subset meets the critical condition $M_{T}\cap M_{I}=\emptyset$, ensuring no overlap in modalities between training and inference.
Within this framework, we define two distinct datasets: one for the training phase and another for the inference phase. The training dataset $D_{T}$ consists of $n_{T}$ samples. Each sample $x$ is restricted to $M_{T}$, denoted as $D_{T}=\{(x_{M_{T}}^{i},y^{i})\}_{i=1}^{n_{T}}$. Similarly, the inference dataset $D_{I}$ consists of $n_{I}$ samples, each restricted to $M_{I}$, formalized as $D_{I}=\{(x_{M_{I}}^{i},y^{i})\}_{i=1}^{n_{I}}$. Our algorithms are designed to build the model $F$ on $D_{T}$ and evaluate unseen data and modality combinations in $D_{I}$. This evaluation measures the model’s ability to generalize knowledge in zero-shot multimodal learning.
@@ 3.2. Text Transformation
For the challenge of the heterogeneity between different modalities, we attempt to map the data of different modalities into a similar semantic space at the raw input stage, integrating modalities with minimal processing. Compared to traditional modality learning approaches that require using embedding features and need to train a new embedding layer for each task or new modality which challenges in zero-shot scenarios. Our approach simplifies this by using pre-trained MLLM or rule-based methods to convert raw inputs into text, enhancing efficiency in cross-modality applications.
For image modality, we utilize the state-of-the-art image caption model, which allows us to take images and answer questions about them. In this stage, the image caption model is prompted to describe images and generate detailed text output, effectively translating visual content into descriptive text. For table modality, we implement the simple text serialization template, which is ”The column name is values,” proposed by TabLLM(Hegselmann et al., 2023). According to TabLLM’s study, this template is proven to outperform zero-shot tabular learning in the LLM. For text modality, we keep human-written sentence forms without conversion to preserve the natural linguistic structure. After transformation, the sentence order of each sample $x$ is determined by the order of modalities in $M$ to concatenate the text sentences and is used as the input for the next step. We have included and compared various SOTA image caption models in Table 5.
A Real-world Example: When predicting diseases, we often have access to patients’ pathology table reports, medical imaging, and audio of patient narration. First, we will perform text transformation on these data. For the images, we use an image caption model that outputs descriptions such as ”The patient has sigmoid colon cancer causing an obstruction, which has led to dilation in the descending colon.” For the tables, we transform data with the template and output statements like ”Histologic Type is Adenocarcinoma” and ”Histologic Grade is Moderately differentiated.” For the audio files, we use the audio caption model that outputs descriptions such as ”I’ve been a little bloated for two weeks, and I have had only three bowel movements.
@@ 3.3. Text-style Translation across Modality
After text transformation, we obtain textual representations for each modality. During the inference phase, we employ LLMs to perform text-style translation across different modalities, reducing the mismatch between training and inference data. This is because, although all types of data are converted into textual representations, there are still syntactic and semantic gaps between the transformed text across different modalities, which are highly dependent on the foundation models used for transformation. For instance, the textual representations of tabular data can follow a static template, while the textual representation of the image is affected by the language style of the image caption model. This divergence in textual representations leads to gaps in semantic space distribution. Our solution is to adopt in-context learning in LLMs to translate the text style to the one in the target space. Text-style translation function $T$ can be formulated as follows:
$T:M_{I}\rightarrow M_{T}$
Specifically, we enhance the capability of the LLMs to effectively interpret the semantic content and convert the modality text from $M_{I}$ format into $M_{T}$ format. This transformation is achieved through few-shot in-context learning, where LLMs are exposed to three examples, which are samples from $D_{T}$, illustrating the desired translation for our modality combination used in training and inference. Moreover, to mitigate information loss during conversion, we specifically emphasize maintaining the integrity of the original information in our prompt. This emphasis ensures that the essential information content and cross-modal connections are preserved during the cross-modal translation process. Figure. 4 demonstrates the details of the Text-style translation prompt.
A Real-world Example: When the training combination for disease prediction includes table modality, and only video modality data is available at inference, we will perform text-style translation on the textual representation of audio data. Continuing the example from Section 3.2, the textual representation of the audio, ”I’ve been a little bloated for two weeks, and I have had only three bowel movements,”, is translated as ”Symptom is Bloating. The symptom is difficulty with bowel movements. Duration is Two weeks.